# Cospectral vertices, Walk-regular planar graphs and the echolocation problem

### v1

@@ 4. Walk-regular planar graphs
($d=3,r_{1}=5,k=1$) Each face with $m\geq 6$ edges must be surrounded by exact $m/2$ pentagonal faces and thus $m$ is even. As $p(v_{1})>\frac{1}{2}$ by (4.1), the only possibility is $r(v_{1})=(5,6,6)$. Suppose $r(v)\not\equiv(5,6,6)$. Then there exists a pentagonal face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$. The $3$-neighborhood of $\{v,v^{\prime}\}$ is unique, with 23 vertices (Figure 6). This gives
$A_{v,v}^{6}=91>89=A^{6}_{v^{\prime},v^{\prime}}$
Hence $r(v)\equiv(5,6,6)$, the graph is isomorphic to the $(5,6,6)$-solid.
Figure 6. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$.
@@
($d=5,r_{1}=3,k=4$) As $p(v_{1})>\frac{3}{2}$ by (4.1), $r_{5}(v_{1})=4$ or $5$. When $r_{5}(v_{1})=4$, we suppose $r_{5}(v)\not\equiv 4$. Then there exist two adjacent vertices $v,v^{\prime}$ such that $r_{5}(v)=4<r_{5}(v^{\prime})$. Counting closed walks of length $4$ in the $2$-neighborhood of $\{v,v^{\prime}\}$ (Figure 7), we have
$A_{v,v}^{4}=59>57=A_{v^{\prime},v^{\prime}}^{4}.$
Hence $r(v)\equiv(3,3,3,3,4)$, the graph is isomorphic to the $(3,3,3,3,4)$-solid.
Figure 7. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,4)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>4$.
In the case that $r_{5}(v_{1})=5$, $r_{5}(v)\not\equiv 5$, there must exist two adjacent vertices $v,v^{\prime}$ with $r_{5}(v)=5<r_{5}(v^{\prime})$. Counting closed walks of length 5 in the 2-neighborhood of $\{v,v^{\prime}\}$ (Figure 8) gives
$A_{v,v}^{5}=182>180=A_{v^{\prime},v^{\prime}}^{5}.$
Hence $r(v)\equiv(3,3,3,3,5)$, the graph is isomorphic to the $(3,3,3,3,5)$-solid.
Figure 8. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,5)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>5$.
We now move on to the more involved cases.
Case (3)
($d=3,r_{1}=3,k=1$) Denote the vertex set of the triangular face $f_{1}(v_{1})$ by $\{v_{1},v_{2},v_{3}\}$, where $v_{2}$ does not belong to the face $f_{2}(v_{1})$. Note that every face with $m\geq 4$ edges should be surrounded by exact $m/2$ triangular faces, and thus $m$ must be even. Suppose $r_{2}(v_{1})>10$. Then
$p(v_{1})\leq\frac{1}{3}+\frac{1}{12}+\frac{1}{12}=\frac{1}{2},$
which violates (4.1). Hence $r_{2}(v_{1})=4,6,8$ or $10$.
Sub-case: $r_{2}(v_{1})=4$
If $r_{3}(v_{1})>4$, it follows that $r_{2}(v_{2})>4$ and
$A_{v_{1},v_{1}}^{4}=17>15=A_{v_{2},v_{2}}^{4}.$
Hence the walk-regularity implies $r_{3}(v_{1})=4$. In this case, the only possibility is $r(v)\equiv(3,4,4)$, and the graph is isomorphic to the triangular prism.
Sub-case: $r_{2}(v_{1})=6$
Suppose $r_{3}(v_{1})>6$. Then $r_{2}(v_{2})>6$, otherwise we have $r(v_{3})=(3,6,6)$ that violates the minimality of $p(v_{1})$. Counting the closed walks with length $6$ in the $3$-neighborhood of $\{v_{1},v_{2}\}$ (Figure 9)
$A_{v_{1},v_{1}}^{6}=91>89=A_{v_{2},v_{2}}^{6}.$
Thus $r(v_{1})=(3,6,6)$ by the walk-regularity.
Figure 9. 3-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,6,m)$ with $m>6$.
Now we claim that $r(v)\equiv(3,6,6)$. Indeed, whenever this is false, there exists a triangular face with vertex set $\{v,v^{\prime},v^{\prime\prime}\}$ such that $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$ (Figure 10), which gives
$A_{v,v}^{6}=93>91=A_{v^{\prime},v^{\prime}}^{6}.$
Hence $r(v)\equiv(3,6,6)$, and the graph is isomorphic to the $(3,6,6)$-solid.
Figure 10. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$.
Sub-case: $r_{2}(v_{1})=8$
As in the previous sub-case, $r_{3}(v_{1})>8$ implies $r_{2}(v_{2})>8$, and a counting in the $4$-neighborhood of the pair $\{v_{1},v_{2}\}$ (Figure 11) provides
$A_{v_{1},v_{1}}^{8}=593>591=A_{v_{2},v_{2}}^{8},$
thus $r(v_{1})=(3,8,8)$.
Figure 11. 4-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,8,m)$ with $m>8$.
If $r(v)\equiv(3,8,8)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$ (Figure 12), and
$A_{v,v}^{8}=595>593=A_{v^{\prime},v^{\prime}}^{8}.$
This shows that $r(v)\equiv(3,8,8)$, and the graph is isomorphic to the $(3,8,8)$-solid.
Figure 12. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$.
Sub-case: $r_{2}(v_{1})=10$
Again, $r_{3}(v_{1})>10$ implies $r_{2}(v_{2})>10$, and a counting in the 5-neighborhood of $\{v_{1},v_{2}\}$ (Figure 13) gives
$A_{v_{1},v_{1}}^{10}=4225>4223=A_{v_{2},v_{2}}^{10},$
thus $r(v_{1})=(3,10,10)$.
Figure 13. 5-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,10,m)$ with $m>10$.
If $r(v)\equiv(3,10,10)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$ (Figure 14), and
$A_{v,v}^{10}=4227>4225=A_{v^{\prime},v^{\prime}}^{10}.$
This shows that $r(v)\equiv(3,10,10)$, and the graph is isomorphic to the $(3,10,10)$-solid.
Figure 14. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$.
Case (6)
($d=3,r_{1}=4,k=1$) Each face with $m\geq 5$ should be surrounded by exact $m/2$ quadrangular faces and thus $m$ is even. As $p(v_{1})>\frac{1}{2}$ by (4.1), we must have $r_{2}(v_{1})=6$ and $r_{3}(v_{1})=6,8$ or $10$.
Sub-case: $r_{3}(v_{1})=6$
Suppose $r(v)\not\equiv(4,6,6)$. Then there exists a quadrangular face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$ (Figure 15). By counting closed walks we obtain
$A_{v,v}^{6}=115>113=A_{v^{\prime},v^{\prime}}^{6}.$
Thus $r(v)\equiv(4,6,6)$, and the graph is isomorphic to the $(4,6,6)$-solid.
Figure 15. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$.
Sub-case: $r_{3}(v_{1})=8$
For this we claim that for every vertex $v$ with $r(v)=(4,6,8)$, its adjacent vertex $v^{\prime}$ outside the hexagonal face $f_{2}(v)$ also has $r(v^{\prime})=(4,6,8)$. Indeed, if this is violated by some adjacent pair $v,v^{\prime}$ (Figure 16), then
$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
This verifies the claim.
Figure 16. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r_{2}(v^{\prime})>6$.
Now suppose $r(v)\not\equiv(4,6,8)$. Then by the claim there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$ (Figure 17). A counting of closed walks with length $8$ gives
$A_{v,v}^{8}=811>809=A_{v^{\prime},v^{\prime}}^{8}.$
Hence $r(v)\equiv(4,6,8)$, and the graph is isomorphic to the $(4,6,8)$-solid.
Figure 17. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$.
Sub-case: $r_{3}(v_{1})=10$
Note that any two adjacent vertices $v,v^{\prime}$ with $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$ (Figure 18) satisfy
$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
This shows that any vertex that is adjacent to some vertex with label $(4,6,10)$ should lie on a hexagonal face.
Figure 18. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$.
Moreover, if $r(v)\not\equiv(4,6,10)$, then there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$ (Figure 19), which give
$A_{v,v}^{10}=6065>6063=A_{v^{\prime},v^{\prime}}^{10}.$
Hence $r(v)\equiv(4,6,10)$, and the graph is isomorphic to the $(4,6,10)$-solid.
Figure 19. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$.
Case (12)
($d=4,r_{1}=3,k=2$) We start with a lemma.
Lemma 4.3.
Let $G$ be $d$-regular with $d\geq 3$. Then,
$A_{v_{i},v_{i}}^{4}-(2d^{2}-d)=2(\#\text{ of $4$-cycles through $v_{i}$}).$
Proof.
Every closed walk of length $4$ is either contractible or traces a $4$-cycle in one of two directions, and there are exactly $d(d-1)+d^{2}$ contractible closed walks of length $4$. ∎
We understand each vertex is adjacent to two triangles and two other faces, say of degree $4\leq m\leq m^{\prime}$ respectively, both larger than $3$. We subdivide the argument into four sub-cases where there are zero, one, two, or more than two $4$-cycles through each vertex.
Sub-case: Zero 4-cycles
There are no squares, and no two triangles share an edge. Such graphs have plenty of structure. Firstly, each face of $\geq 5$ sides is adjacent only to triangular faces. Furthermore, each such face is, through each vertex, opposed to another face of $\geq 5$ sides. We first establish that, in this situation,
$A_{v_{i},v_{i}}^{5}-60=2\times(\#\text{ of pentagons incident to $v_{i}$})$
Our job is now to eliminate the cases where
$r(v)=(3,5,3,6)\text{ or }(3,5,3,7)$
for each $v$, and we will be left with $r(v)\equiv(3,5,3,5)$. In fact, if we have either case above for one vertex, we have it for all of them. We can construct a $3$-neighborhood of the vertex $v$ (Figure 20) with either a hexagon or a face of $\geq 7$ sides and obtain
$A_{v,v}^{6}=270+\begin{cases}2&r(v)=(3,5,3,6)\\ 0&r(v)=(3,5,3,7).\end{cases}$
Figure 20. Two possible 3-neighborhood of $v$.
If we attempt to construct either graph with either $r\equiv(3,5,3,6)$ or $r\equiv(3,5,3,7)$, we quickly run into absurdities. Hence, we are left with the $(3,5,3,5)$ solid.
Sub-case: one $4$-cycle
@@
Note, one or both of the antennae may attach to a wing to form a pentagon. Running some software, we find that, for any vertex $v$ of the square,
$A_{v,v}^{5}=\begin{cases}68&\text{ if no pentagon is incident to $v$}\\ 70&\text{ if the remaining face incident to $v$ is a pentagon}.\end{cases}$
On the other hand, the vertex $v^{\prime}$ at the “neck” of the bee satisfies
$A_{v^{\prime},v^{\prime}}^{5}=62+2\times(\#\text{ of pentagons incident to $v^{\prime}$})\leq 66,$
and so we have $A_{v^{\prime},v^{\prime}}^{5}<A_{v,v}^{5}$, a contradiction.
We have shown that the only $4$-cycles we find are the boundary of the union of two triangular faces. Since there are no squares, there must be at least one pentagon. The graph self-destructs upon any attempt to draw a $1$-neighborhood of the pentagon.
Sub-case: two $4$-cycles
There are a few situations we must rule out. We first establish that each vertex must be incident to a square face. Suppose $v$ is not adjacent to a square face. Then, $r(v)=(3,3,m,m^{\prime})$ or $(3,m,3,m^{\prime})$, where $m,m^{\prime}\geq 5$. The first is incident to only one $4$-cycle, so is ruled out. We try to draw the graph starting at this vertex and obtain the one in Figure 22, which forces a vertex to have at least three $4$-cycles.
@@
Sub-case: More than two $4$-cycles
@@
Hence, all that remains is to exclude the case where the vertices are incident to exactly three $4$-cycles.
A similar argument as the one above excludes the case where $v$ is incident to a face of $5$ or more sides. Hence, for each vertex $v$, we have either $r(v)=(3,4,3,4)$ or $r(v)=(3,3,4,4)$. There is a unique graph with these properties, depicted in Figure 24.
@@
This graph, however, is not walk-regular. The three innermost and outermost vertices all have $74$ returning walks of length $5$, while the other vertices have $76$.
@@
($d=4,r_{1}=3,k=1$) As in the proof of the previous case, we subdivide the argument according to the number of 4-cycles through each vertex. Since there is exactly 1 triangle face adjacent to each vertex, the number of 4-cycles is determined by the number of adjacent squares. Therefore, each vertex must be adjacent to the same number of squares. By (4.1), we must have $p(v_{1})>1$, and thus there must be at least two squares adjacent to each vertex.
Sub-case: three squares at each vertex
This case is straightforward since we already know that $r(v)\equiv(3,4,4,4)$. We remark that, in this case, when one wants to reconstruct the $(3,4,4,4)$-solid, there are two possible graphs, the other one being the twisted $(3,4,4,4)$-solid graph. However, there are vertices on the twisted $(3,4,4,4)$ solid that have different numbers of closed walks of length 7, and thus, this twisted case should be ruled out. See Figure 25.
@@
Sub-case: two squares at each vertex
In this case, $r(v_{1})$ is either $(3,4,4,5)$ or $(3,4,5,4)$. We first consider the case $r(v_{1})=(3,4,4,5)$. One can try to draw such a graph starting from $v_{1}$ and end up with a neighborhood of the shape as in Figure 26.
@@
Here, vertex $v_{1}$ is what we started with. We claim that the unknown polygon adjacent to vertex $v_{3}$ must be a pentagon. Then a similar argument will show that all outer polygons must be pentagons. To do so, we count the closed walks of length 5 for $v_{1}$ and for $v_{3}$. Non-trivial ones are either a pentagon or a triangle attached to a square, which we shall call a “house”. $v_{1}$ and $v_{3}$ always share the same number of “houses”, and we are done with the claim.
Next, we claim that we will get a contradiction if the center polygon is not a pentagon. In this case, for $v_{1}$ and $v_{2}$ to have the same number of closed walks of length 5, we cannot create more “houses” in the outer layer for vertex $v_{1}$. This forces us to get 3 squares in a row. After drawing all the squares we arrive at Figure 27, which is impossible since there are vertices adjacent to more than 2 triangles.
Figure 27. An impossible graph when the center is not a pentagon.
Therefore, we conclude that we must have a pentagon at the center, and this in turn implies that we must add an additional “house” to vertex $v_{1}$, so we arrive at Figure 28.

### v2

@@ 4.2. The cubic cases
($d=3,r_{1}=5,k=1$) Using (4.1) we immediately get $\ell=r_{2}(v_{1})=6$. We will prove that every $6$-cycle is a face in this case, and conclude the proof of Lemma 4.6.
Proof of Lemma 4.6 for $r_{1}=5$.
Suppose that a 6-cycle $C=\overline{a_{1}a_{2}\cdots a_{6}a_{1}}$ is not a face. Without loss of generality, we further assume that such $C$ is minimal among all the 6-cycles that are not faces. More precisely, this minimality of $C$ implies that any 6-cycle properly contained in the closed region bounded by $C$ should be a face.
By the 3-connectedness of the graph, $C$ has exactly three inner incident edges and three outer incident edges. There is a unique (up to isomorphism) possible neighborhood of $C$ such that $\mathcal{C}_{5}(a_{i})=1$ for all vertices $a_{i}$ on $C$, as depicted in Figure 6, where $C$ is adjacent to a face $\overline{b_{2}a_{4}a_{5}a_{6}b_{3}b_{2}}$ with length 5 inside $C$ and a pentagon face outside $C$ bounded by the path $\overline{a_{1}a_{2}a_{3}}$. Note that the inner neighbor $b_{1}$ of $a_{2}$ is not adjacent to $b_{2}$ or $b_{3}$, and hence, both the face bounded by the path $\overline{b_{1}a_{2}a_{1}a_{6}b_{3}}$ and the one bounded by $\overline{b_{1}a_{2}a_{3}a_{4}b_{2}}$ have length $>5$. Besides these two faces, the remaining adjacent face of $b_{1}$ should be of length $5$ so that $\mathcal{C}_{5}(b_{1})=1$. However, no matter which adjacent face of $b_{1}$ is of length $6$, the minimality of $C$ implies that $\mathcal{C}_{6}(b_{1})<\mathcal{C}_{6}(a_{2})$, which contradicts Lemma 4.1.
Figure 6. An impossible 6-cycle with $r_{1}=5$.
Now using Lemmas 4.1 and 4.6 we have $\mathcal{C}_{6}(v)\equiv\mathcal{C}_{6}=1$ or $2$. A pentagon face on which each vertex is adjacent to exactly one face with length 6 is simply impossible, and hence $\mathcal{C}_{6}=2$. This yields that $r(v)\equiv(5,6,6)$, and the graph is isomorphic to the $(5,6,6)$-solid.
4.3. The 4-regular cases
Now we assume that $G$ is 4-regular. We first justify that every triangle in $G$ is a face:
Proof of Lemma 4.3 for $d=4$.
Suppose that a triangle $C$ is not a face. By the 3-connectedness, each vertex $a_{i}$ has at least one incident edge inside $C$ and at least one incident edge outside $C$. Therefore, it must have $d_{*}(C)=(1,1,1)$. However, this is simply impossible as the total degree of the induced subgraph of the vertices inside $C$ is odd. ∎
Consequently, the octahedron is determined as $r(v)\equiv(3,3,3,3)$ in case (10). For case (11) ($d=4,r_{1}=3,k=3$), the face $f_{4}(v_{1})$ surrounded by $m=r_{4}(v_{1})\geq 4$ triangular faces determines the unique pattern up to isomorphism, hence $r(v)\equiv(3,3,3,m)$ and the graph is isomorphic to the $m$-gonal antiprism.
The other two cases (12) and (13) are more involved. We need to prove that every 4-cycle is either a face or the boundary of two adjacent triangular faces there.
Proof of Lemma 4.4 for $d=4$.
Suppose that a chordless 4-cycle $C=\overline{a_{1}a_{2}a_{3}a_{4}a_{1}}$ is not a face. By the 3-connectedness, at least three of $\{a_{1},a_{2},a_{3},a_{4}\}$ have incident edges inside $C$, and at least three of them have incident edges outside $C$. Without loss of generality, we assume that such $C$ is minimal, that is, any chordless 4-cycle properly contained in the closed region bounded by $C$ should be a face.
Since the total degree of the subgraph inside $C$ is even, the number of incident edges of $\{a_{1},a_{2},a_{3},a_{4}\}$ inside $C$ equals $4$. The possible inner degree sequence of $C$ is $(0,1,1,2),(0,1,2,1)$ or $(1,1,1,1)$. The case $d_{*}(C)=(0,1,2,1)$ is simply ruled out, as $\{a_{2},a_{4}\}$ should not be a cut-set of the graph.
Case: $d_{*}(C)=(0,1,1,2)$. Note that $\mathcal{C}_{3}=\mathcal{C}_{3}(v_{1})\geq 1$. In this case, $\mathcal{C}_{3}\geq 2$ is impossible since $\{a_{1},a_{2},a_{3},a_{4}\}$ has $\geq 3$ neighbors inside $C$ and $\geq 3$ neighbors outside $C$ (by the 3-connectedness). It follows that $\mathcal{C}_{3}=1$, and every 4-cycle inside $C$ is chordless so is a face. Using (4.1), we have $\mathcal{C}_{4}=\mathcal{C}_{4}(v_{1})\geq 2$. Lemma 4.3 implies that every vertex inside $C$ is surrounded by at most one face with length $>4$. Let $b_{1},b_{2}$ be the two neighbors of $a_{4}$ inside $C$. When $\overline{b_{1}b_{2}}$ is an edge, to have $\mathcal{C}_{3}(a_{3})=\mathcal{C}_{3}(a_{4})=1$, $a_{2}$ and $a_{3}$ have exactly one common neighbor $u$, and this gives a cut-set $\{a_{4},u\}$ if $u$ lies inside $C$ or a cut-set $\{a_{1},u\}$ if $u$ lies outside $C$, which violates the 3-connectedness.
Now the only possibility left is that one of $b_{1},b_{2}$ (say $b_{1}$) is adjacent to $a_{3}$, and one of the neighbors of $a_{1}$ outside $C$ is adjacent to $a_{2}$. Let $b_{3}$ be the neighbor of $a_{2}$ inside $C$. Since the face bounded by $\overline{b_{2}a_{4}a_{1}a_{2}b_{3}}$ has length $>4$, we have $\mathcal{C}_{4}=2$, and both the face bounded by $\overline{b_{2}a_{4}b_{1}}$ and the one bounded by $\overline{b_{3}a_{2}a_{3}b_{1}}$ have length $4$, as drawn in Figure 7. Then, to have $\mathcal{C}_{3}(b_{1})+\mathcal{C}_{4}(b_{1})=3$, the undetermined face with $\overline{b_{1}b_{3}}$ should have length $>4$, contradicting the fact that $\mathcal{C}_{3}(b_{3})+\mathcal{C}_{4}(b_{3})=3$.
Figure 7. An impossible 4-cycle with $d_{*}(C)=(0,1,1,2)$.
Case: $d_{*}(C)=(1,1,1,1)$. One special situation is that $\{a_{1},a_{2},a_{3},a_{4}\}$ has a common neighbor inside $C$, in which $\mathcal{C}_{3}=4$ and the graph is isomorphic to the octahedron. If this is not the case, again by the 3-connectedness, there are at least three distinct neighbors of $\{a_{1},a_{2},a_{3},a_{4}\}$ inside $C$ as well as at least three distinct ones outside $C$. From this we know that $\mathcal{C}_{3}=1$, one edge (say $\overline{a_{1}a_{4}}$) of $C$ lies on an incident triangle inside $C$, and the opposite edge ($\overline{a_{2}a_{3}}$) of $C$ lies on an incident triangle outside $C$. Denote by $b_{1},b_{2},b_{3}$ the neighbors of $a_{1},a_{2},a_{3}$ inside $C$ respectively. Note that there are no chorded $4$-cycles inside $C$ or else $\mathcal{C}_{3}\geq 2$, and hence every $4$-cycle inside $C$ is a face by the minimality of $C$, thus $\mathcal{C}_{4}\leq 3$.
If $\mathcal{C}_{4}=3$, then none of the faces inside $C$ has length $>4$, and it forces $\overline{b_{1}b_{2}b_{3}b_{1}}$ to be a triangle but not a face, which violates Lemma 4.3.
Now we suppose $\mathcal{C}_{4}=2$. Once $b_{2},b_{3}$ are adjacent, to have $\mathcal{C}_{4}(a_{2})=\mathcal{C}_{4}(a_{3})=2$, both the face bounded by $\overline{b_{2}a_{2}a_{1}b_{1}}$ and the one bounded by $\overline{b_{3}a_{3}a_{4}b_{1}}$ have length $>4$, which yields that $\mathcal{C}_{4}(b_{1})\leq 1$. Hence any graph with $\overline{b_{2}b_{3}}$ is ruled out.
Without $\overline{b_{2}b_{3}}$, the face bounded by $\overline{b_{2}a_{2}a_{3}b_{3}}$ has length $>4$. To have $\mathcal{C}_{4}(b_{2})=\mathcal{C}_{4}(b_{3})=2$, both $b_{2},b_{3}$ must be adjacent to $b_{1}$, which gives a cut-set $\{b_{2},b_{3}\}$ of the graph (Figure 8), a contradiction.
Figure 8. An impossible 4-cycle with $d_{*}(C)=(1,1,1,1)$.
@@ 4.3. The 4-regular cases
($d=4,r_{1}=3,k=2$) We first justify Lemma 4.5 for this case.
Proof of Lemma 4.5 for case (12).
Suppose a chordless 5-cycle $C=\overline{a_{1}a_{2}a_{3}a_{4}a_{5}a_{1}}$ is not a face. Note that the number of inner incident edges of $C$ must be even. By the 3-connectedness, the possible choices for $d_{*}(C)$ are $(0,0,1,1,2)$, $(0,1,0,1,2)$, $(0,1,1,1,1)$, $(0,1,1,2,2)$, $(0,1,2,1,2)$ or $(1,1,1,1,2)$. In view of the symmetry between the inside and outside of $C$, it suffices to rule out the first three situations, as illustrated in Figure 9.
Figure 9. Three impossible 5-cycles with $d_{*}(C)=(0,0,1,1,2)$, $(0,1,0,1,2)$ and $(0,1,1,1,1)$ respectively.
Sub-case: $d_{*}(C)=(0,0,1,1,2)$. Note that one of the two incident triangles of $a_{1}$ is bounded by $\overline{a_{1}a_{2}}$, and one of the two incident triangles of $a_{5}$ is bounded by $\overline{a_{4}a_{5}}$. To have three distinct inner neighbors of $C$ (by the 3-connectedness), $a_{3}$ has no incident triangles inside $C$, and thus $a_{2},a_{3},a_{4}$ have a common outer neighbor. This leads to $\mathcal{C}_{4}(a_{4})\geq 2>\mathcal{C}_{4}(a_{5})$, a contradiction.
Sub-case: $d_{*}(C)=(0,1,0,1,2)$. As in the first sub-case, a triangle outside $C$ is bounded by $\overline{a_{1}a_{2}}$ and one inside $C$ is bounded by $\overline{a_{4}a_{5}}$. It follows similarly that $a_{2}$ has no inner incident triangles, and this forces $a_{1},a_{2},a_{3}$ to have a common neighbor outside $C$ which has three incident triangles.
Sub-case: $d_{*}(C)=(0,1,1,1,1)$. By the 3-connectedness, there is at most one inner incident triangle of $C$. We also note that, in $\{a_{2},a_{3},a_{4},a_{5}\}$ the vertices with two incident triangles outside $C$ cannot be adjacent, otherwise four of the vertices on $C$ have a common neighbor that is incident to three triangles. Therefore, the only possible choice of the unique inner incident triangle of $C$ is the one bounded by $\overline{a_{3}a_{4}}$. However, in this situation $C$ has only two outer neighbors, which violates the 3-connectedness. ∎
Now we are ready to finish off Case (12). In this case, each vertex is incident to two triangles and two other faces of length $\geq 4$. In view of Lemma 4.1, we subdivide the argument into four sub-cases where $\mathcal{C}_{4}(v)\equiv\mathcal{C}_{4}=0,1,2$ or $\geq 3$.
Sub-case: $\mathcal{C}_{4}=0$
There are no faces of length $4$, and no two triangles share an edge. It follows that for each $v$, $r(v)$ is of the form $(3,m,3,m^{\prime})$ for some $5\leq m\leq m^{\prime}$. Such graphs have plenty of structure. Firstly, each face of length $\geq 5$ is adjacent only to triangular faces. Furthermore, each such face is, through each vertex, opposed to another face of length $\geq 5$.
Using Lemma 4.1 and (4.1), we have $\mathcal{C}_{5}(v)=\mathcal{C}_{5}(v_{1})\geq 1$ for any vertex $v$. Also we note that every 5-cycle is chordless, and so is a face by Lemma 4.5. If we attempt to construct a triangle on which each vertex is incident to exactly one pentagon face, we quickly run into absurdities. Hence, we are left with $r(v)\equiv(3,5,3,5)$, and the graph is isomorphic to the $(3,5,3,5)$-solid.
Sub-case: $\mathcal{C}_{4}=1$
@@
Note, one or both of the antennae may attach to a wing to form a pentagon face. Applying Lemma 4.5, we find that, $\mathcal{C}_{5}(v)\geq 4$ for any vertex $v$ on the square face, while $\mathcal{C}_{5}(v^{\prime})\leq 3$ for the vertex $v^{\prime}$ at the “neck” of the bee. This is contrary to Lemma 4.1.
Combining this and Lemma 4.4, we have shown that the only $4$-cycles we find are the boundary of the union of two triangular faces. Since there are no square faces, there must be at least one pentagon face. The graph self-destructs upon any attempt to draw the $1$-neighborhood of a pentagon face.
Sub-case: $\mathcal{C}_{4}=2$
There are a few situations we must rule out. We first establish that each vertex must be incident to a square face. Suppose $v$ is not adjacent to a square face. Then, $r(v)=(3,3,m,m^{\prime})$ or $(3,m,3,m^{\prime})$, where $m,m^{\prime}\geq 5$. The first is incident to only one $4$-cycle, so is ruled out. We try to draw the graph starting at this vertex and obtain the one in Figure 11, which forces a vertex to lie on at least three 4-cycles.
@@
Sub-case: $\mathcal{C}_{4}\geq 3$
@@
Hence, all that remains is to exclude the case $\mathcal{C}_{4}=3$.
A similar argument as the one above excludes the case where $v$ is incident to a face of length $\geq 5$. Hence, for each vertex $v$, we have either $r(v)=(3,4,3,4)$ or $r(v)=(3,3,4,4)$. There is a unique graph with these properties, depicted in Figure 13.
@@
This graph, however, is not walk-regular. The three innermost and outermost vertices all have seven incident $5$-cycles, while the other vertices have eight.
@@
($d=4,r_{1}=3,k=1$) As in the proof of the previous case, we subdivide the argument according to $\mathcal{C}_{4}$, the number of 4-cycles through each vertex. Since there is exactly 1 triangular face incident to each vertex, $\mathcal{C}_{4}$ is exactly the number of square faces around it. By (4.1), we must have $p(v_{1})>1$, and thus $\mathcal{C}_{4}\geq 2$.
Sub-case: $\mathcal{C}_{4}=3$
This case is straightforward since we already know that $r(v)\equiv(3,4,4,4)$. We remark that, however, when one wants to reconstruct the $(3,4,4,4)$-solid, there are two possible graphs, the other one being the twisted $(3,4,4,4)$-solid graph. Nonetheless, there are vertices on the twisted $(3,4,4,4)$-solid that have different numbers of closed walks of length 7, and thus, this twisted case should be ruled out. See Figure 14.
@@
Sub-case: $\mathcal{C}_{4}=2$
In this sub-case, $r(v_{1})$ is either $(3,4,4,5)$ or $(3,4,5,4)$. We first consider the case $r(v_{1})=(3,4,4,5)$. One can try to draw such a graph starting from $v_{1}$ and end up with a neighborhood of the shape as in Figure 15.
@@
Here, vertex $v_{1}$ is what we started with. We claim that the unknown face incident to vertex $v_{3}$ must be of length 5. Then a similar argument will show that all unknown faces within the 2-neighborhood of the center face must be of length 5. To do so, we count the 5-cycles incident to $v_{1}$ and to $v_{3}$. Note that $v_{1}$ and $v_{3}$ always share the same number of “houses” (i.e., the 5-cycle with a chord), and $v_{1}$ is already incident to a pentagon face. To satisfy $\mathcal{C}_{5}(v_{1})=\mathcal{C}_{5}(v_{3})$, we must have a chordless 5-cycle $C$ through $v_{3}$, and there are three possible situations: (i) $C=\overline{u_{1}u_{2}v_{3}u_{3}u_{4}u_{1}}$, (ii) $C=\overline{u_{2}v_{3}u_{3}u_{4}u_{5}u_{2}}$, and (iii) $u_{2},u_{4}$ have a common neighbor $w$ such that $C=\overline{wu_{2}v_{3}u_{3}u_{4}w}$.
Among these situations, (i) is simply ruled out as $\{u_{2},u_{4}\}$ would be a cut-set of the graph. For (ii), the 3-connectedness yields that each of $u_{2}$ and $u_{5}$ has an incident edge inside $C$, and these two edges must have a common endpoint $x$ to bound the incident triangle of $u_{2}$. However, there is still a cut-set $\{x,u_{4}\}$ of the graph. Now we are left with (iii). Using a similar argument, it is impossible to have $\geq 3$ inner neighbors of $C=\overline{wu_{2}v_{3}u_{3}u_{4}w}$, which is a contradiction. Hence, this $C$ must be a face with length $5$, and thus the claim is justified.
Next, we claim that we will get a contradiction if the center face is of length $\geq 6$. In this case, to have $\mathcal{C}_{5}(v_{1})=\mathcal{C}_{5}(v_{2})$, we cannot create more “houses” in the outer layer for vertex $v_{1}$. This forces us to get three square faces in a row. After drawing all the square faces we arrive at Figure 16, which is impossible since there are vertices incident to more than 2 triangles.
Figure 16. An impossible graph when the center face is of length $\geq 6$.
Therefore, we conclude that we must have a pentagon face at the center, and this in turn implies that we must add an additional “house” to vertex $v_{1}$, so we arrive at Figure 17.